# Policy Contrastive Decoding for Robotic Foundation Models

### v2

@@ 5 Conclusion, Limitations and Future Work
While our PCD demonstrates effectiveness and flexibility, there are certain limitations. 1) PCD performs object detection and segmentation for each visual observation in the trajectory to produce object-masked observations, which increases the computational overhead of the baseline police. Recent progress in fast LLM inference techniques [55, 23, 30] could potentially enhance the computational efficiency of our PCD approach. 2) This work concentrates exclusively on addressing the spurious correlation issue in robot policies during the testing stage, without exploring methods to prevent the learning of such correlations during training. Hence, future work will focus on investigating effective strategies to tackle spurious correlations at both the training and inference levels.
Appendix A Appendix
A.1 CAM Visualization of Failure Cases.
We present the CAM [38] visualization of failure cases for the OpenVLA [^1] We only visualize the CAM results of OpenVLA, as the action prediction mechanism of diffusion-based models (e.g., Octo and $\pi_{0}$) makes it difficult to produce their CAM results. model across the nine tasks in Fig. 6. For each task, the averaged CAM results over the seven action dimensions ([$\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper]) are plotted. As can be observed from the figure, the model tends to predict actions based on spurious or background features in visual observations. For instance, on the “move the blue plastic bottle near sponge” (denoted as “Move Near”) task, the model model habitually focuses on the “coke can” object that it frequently encountered during the training phase. These observations demonstrate that the learned policy utilizes perceptual data for decision-making in a way that significantly diverges from human strategies. Robot policies often erroneously associate task-irrelevant features with actions, a major cause of overfitting and poor generalization across tasks.
@@ Appendix A Appendix
A.2 Details of the Track2Mask Module
In our devised Track2Mask module, the target object specified by the language instruction in the initial observation can be annotated using Point and Box prompts, along with off-the-shelf open-vocabulary object detection models, as illustrated in Fig. 7. The SAM2 [21] model is then applied to automatically track and segment the target object across subsequent observations in the trajectory, which enables precise object masking with minimal or no human intervention during inference. Note that, once the task-specific objects are segmented by the SAM2 model, they are inpainted to create object-masked observations.
@@ A.2 Details of the Track2Mask Module
A.3 Details of the Experimental Setup
A.3.1 Baseline Policies
Simulation Experiments. We conduct simulation experiments using three diverse robot policies in the SIMPLER environment, including the autoregressive policy OpenVLA [20], and diffusion-based policies Octo (base) [32] and $\pi_{0}$ [3].
  • OpenVLA: a vision-language-action model with 7 billion parameters, is trained on 970,000 episodes of robotic demonstrations from the Open X-Embodiment dataset. This policy is fine-tuned using the pre-trained Prismatic [19] model. [^2] https://openvla.github.io/
@@ A.3.1 Baseline Policies
Real-world Experiments. Since the simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA, we conduct real-world experiments using the $\pi_{0}$ policy  [^5] https://www.physicalintelligence.company/blog/pi0. The policy is finetuned using on six real-world tasks, with each task consisting of 10 trajectories. We employ the parameter-efficient finetuning strategy LoRA [16] to adapt $\pi_{0}$ to these real world tasks. The training hyperparameters are listed in Table 2.
@@ A.3.2 Evaluation Tasks
Simulation Close Drawer close top/middle/bottom drawer top/middle/bottom drawer 100
Move Near move X near Y apple, orange, pepsi can, … 100
Open Drawer open top/middle/bottom drawer top/middle/bottom drawer 100
Pick Coke Can pick coke can coke can 100
Apple Drawer open top drawer/place apple into top drawer top drawer, apple 100
Carrot Plate put carrot on plate carrot, plate 100
Eggplant Basket put eggplant into yellow basket eggplant, yellow basket 100
Spoon Towel put the spoon on the towel spoon, towel 100
Stack Cube stack the green block on the yellow block green cube, yellow cube 100
Real-world Pick Ball pick the ball ball 20
@@
A.4 Performance at the Maximum Step
In Table 1, tasks completed before the predefined maximum step are included in the success rate calculation, and the task is terminated upon completion. Table 4 presents success rates computed at the maximum step, irrespective of whether tasks were completed earlier. As seen, PCD demonstrates its advantages by improving the success rates of OpenVLA, Octo and $\pi_{0}$ by up to 33%, 14% and 7%, respectively. We also observe that the vast majority of models underperform on these tasks relative to their results in Table 1, which indicates that even though they complete the task successfully ahead of the maximum steps, their later predicted actions lead to task failure. The superiority of the proposed PCD method is compellingly evidenced by the consistent performance gains over baseline policies presented in Table 1 and Table 4.
Table 4: Success rates (%) of three baseline policies integrated w/ or w/o our PCD over 9 tasks from Google Robot and WidowX—calculated at the maximum step.
OpenVLA[20] Octo [32] $\pi_{0}$ [3]
@@ A.4 Performance at the Maximum Step
Average 0.15 0.20 0.20 0.20 0.14 0.16 0.15 0.16 0.60 0.64 0.64 0.63
+33% +33% +33% +14% +7% +14% +7% +7% +5%
Google Robot Close Drawer 0.41 0.65 0.66 0.76 0.31 0.27 0.37 0.16 0.75 0.71 0.76 0.70
Move Near 0.46 0.46 0.43 0.33 0.05 0.10 0.09 0.12 0.66 0.64 0.62 0.66
Open Drawer 0.20 0.27 0.27 0.23 0.00 0.01 0.00 0.02 0.34 0.44 0.47 0.48
Pick Coke Can 0.24 0.37 0.41 0.37 0.23 0.47 0.38 0.46 0.80 0.86 0.83 0.84
Apple Drawer 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.29 0.24 0.22
WidX. Carrot Plate 0.00 0.02 0.00 0.03 0.10 0.18 0.16 0.12 0.56 0.62 0.57 0.56
Eggplant Basket 0.01 0.01 0.02 0.08 0.47 0.32 0.28 0.40 0.91 0.80 0.80 0.83
Spoon Towel 0.00 0.01 0.00 0.01 0.09 0.08 0.09 0.08 0.81 0.87 0.86 0.83
Stack Cube 0.00 0.01 0.00 0.00 0.00 0.03 0.02 0.05 0.48 0.57 0.59 0.57

### v3

@@ 5 Conclusion, Limitations and Future Work
While our PCD demonstrates effectiveness and flexibility, there are certain limitations. 1) PCD performs object detection and segmentation for each visual observation in the trajectory to produce object-masked observations, which increases the computational overhead of the baseline police. Recent progress in fast LLM inference techniques (Zhou et al., 2024; Leviathan et al., 2023; Liu et al., 2023c) could potentially enhance the computational efficiency of our PCD approach. 2) This work concentrates exclusively on addressing the spurious correlation issue in robot policies during the testing stage, without exploring methods to prevent the learning of such correlations during training. Hence, future work will focus on investigating effective strategies to tackle spurious correlations at both the training and inference levels.
Appendix A CAM Visualization of Failure Cases.
We present the CAM [Selvaraju et al., 2017] visualization of failure cases for the OpenVLA [^1] We only visualize the CAM results of OpenVLA, as the action prediction mechanism of diffusion-based models (e.g., Octo and $\pi_{0}$) makes it difficult to produce their CAM results. model across the nine tasks in Fig. 6. For each task, the averaged CAM results over the seven action dimensions ($\Delta x$, $\Delta y$, $\Delta z$, rotx, roty, rotz, gripper]) are plotted. As can be observed from the figure, the model tends to predict actions based on spurious or background features in visual observations. For instance, on the “move the blue plastic bottle near sponge” (denoted as “Move Near”) task, the model model habitually focuses on the “coke can” object that it frequently encountered during the training phase. These observations demonstrate that the learned policy utilizes perceptual data for decision-making in a way that significantly diverges from human strategies. Robot policies often erroneously associate task-irrelevant features with actions, a major cause of overfitting and poor generalization across tasks.
@@
Appendix B Details of the Track2Mask Module
In our devised Track2Mask module, the target object specified by the language instruction in the initial observation can be annotated using Point and Box prompts, along with off-the-shelf open-vocabulary object detection models, as illustrated in Fig. 7. The SAM2 [Kirillov et al., 2023] model is then applied to automatically track and segment the target object across subsequent observations in the trajectory, which enables precise object masking with minimal or no human intervention during inference. Note that, once the task-specific objects are segmented by the SAM2 model, they are inpainted to create object-masked observations.
@@ Appendix B Details of the Track2Mask Module
Appendix C Details of the Experimental Setup
C.1 Baseline Policies
Simulation Experiments. We conduct simulation experiments using three diverse robot policies in the SIMPLER environment, including the autoregressive policy OpenVLA [Kim et al., 2024], and diffusion-based policies Octo (base) [M. et al., 2024] and $\pi_{0}$ [Black et al., 2024].
  • OpenVLA: a vision-language-action model with 7 billion parameters, is trained on 970,000 episodes of robotic demonstrations from the Open X-Embodiment dataset. This policy is fine-tuned using the pre-trained Prismatic [Karamcheti et al., 2024b] model. [^2] https://openvla.github.io/
@@ C.1 Baseline Policies
Real-world Experiments. Since the simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA, we conduct real-world experiments using the $\pi_{0}$ policy  [^5] https://www.physicalintelligence.company/blog/pi0. The policy is finetuned using on six real-world tasks, with each task consisting of 10 trajectories. We employ the parameter-efficient finetuning strategy LoRA [Hu et al., 2022] to adapt $\pi_{0}$ to these real world tasks. The training hyperparameters are listed in Table 2.
@@ C.2 Evaluation Tasks
Simulation
Close Drawer close top/middle/bottom drawer top/middle/bottom drawer 300
Move Near move X near Y
apple, orange, pepsi can, …
300
Open Drawer open top/middle/bottom drawer top/middle/bottom drawer 300
Pick Coke Can pick coke can coke can 300
Apple Drawer open top drawer/place apple into top drawer top drawer, apple 300
Carrot Plate put carrot on plate carrot, plate 300
Eggplant Basket put eggplant into yellow basket eggplant, yellow basket 300
Spoon Towel put the spoon on the towel spoon, towel 300
Stack Cube stack the green block on the yellow block green cube, yellow cube 300
Real-world
Pick Ball pick the ball ball 20
@@
Appendix D Performance at the Maximum Step
In Table 1, tasks completed before the predefined maximum step are included in the success rate calculation, and the task is terminated upon completion. Table 4 presents success rates computed at the maximum step, irrespective of whether tasks were completed earlier. As seen, PCD demonstrates its advantages by improving the success rates of OpenVLA, Octo and $\pi_{0}$ by up to 45.0%, 19.3% and 7.9%, respectively. We also observe that the vast majority of models underperform on these tasks relative to their results in Table 1, which indicates that even though they complete the task successfully ahead of the maximum steps, their later predicted actions lead to task failure. The superiority of the proposed PCD method is compellingly evidenced by the consistent performance gains over baseline policies presented in Table 1 and Table 4.
Table 4: Success rates (%) and 95% confidence interval of three baseline policies integrated w/ or w/o our PCD over 9 tasks from Google Robot and WidowX—calculated at the maximum step.
OpenVLA[Kim et al., 2024] Octo [M. et al., 2024] $\pi_{0}$ [Black et al., 2024]
@@ Appendix D Performance at the Maximum Step
[…]
+39.1% +39.7% +45.0% +12.6% +17.6% +19.3% +5.9% +7.6% +7.9%
Google
Robot
[…]
Appendix E Computational Overhead
Table 5 reports the computational overhead of three baseline policies integrated w/ or w/o our PCD method in the SIMPLER environment. 1) Inference Latency. PCD approximately doubles the inference latency per step across all three policies. This is an expected trade-off, as PCD requires a second forward pass on the object-masked input to compute the contrastive signal. For Octo and $\pi_{0}$, although the KDE-PM process introduces additional inference time, the overall time overhead is still kept to roughly a twofold increase due to their parallel processing of the original and object-masked images. Crucially, this increased inference time does not significantly affect the total task completion time in real-world scenarios, given the much longer duration of the robot’s physical execution. As proven by Figure 3, PCD only brings 24% extra execution time on OpenVLA. 2) Memory Cost. The impact on memory cost varies based on the policy’s architecture. For OpenVLA, which processes the original and masked inputs serially, the memory overhead is negligible, with the primary cost being the increased sequential processing time. In contrast, the memory costs of Octo and $\pi_{0}$ are $1.22\times$ and $1.37\times$ that of the baseline respectively, due to their parallel data processing mechanisms.
Table 5: Computational overhead of three baseline policies integrated w/ or w/o our PCD method in the SIMPLER environment.
Model OpenVLA Octo $\pi_{0}$
Base +PCD Base +PCD Base +PCD
Average time cost for each infer. step (s) 0.86 1.77 0.21 0.39 0.66 1.09
Memory cost (MB) 16357 16869 2884 3528 8535 11699
Appendix F LLMs Usage Statement
In the preparation of this paper, we employed Large Language Models (LLMs) solely as a writing assistance tool for limited text polishing and language refinement. LLMs were not involved in any aspects of research ideation, conceptual development, technical analysis, algorithm design, experimental execution, or result interpretation. All scientific contributions, methodological innovations, and intellectual content remain entirely our own.